Neelon Corporation has two divisions: Southern Division and Northern Division. The following data are for the most recent operating period: Total Company Southern Division Northern Division Sales $ 334,700 $ 194,900 $ 139,800 Variable expenses $ 121,298 $ 77,960 $ 43,338 Traceable fixed expenses $ 167,200 $ 75,900 $ 91,300 Common fixed expense $ 66,940 $ 38,980 $ 27,960 The common fixed expenses have been allocated to the divisions on the basis of sales. The Northern Division’s break-even sales is closest to:
Answer #1

Contribution Margin for northern division = Sales - Variable costs

= 139,800 - 43,338

= 96,462

Contribution margin ratio = 96,462/139,800 = 69%

Northern Division break even point = Traceable fixed expenses / contribution margin ratio

= 91,300 / 69%

= 132,318.84